How many grams of sodium will result from the decomposition of 1.0 mole of sodium azide


1
Expert's answer
2021-05-19T02:27:18-0400

Solution:

sodium azide = NaN3

Balanced chemical equation:

2NaN3(s) → 2Na(s) + 3N2(g)

According to the equation above: n(NaN3) = n(Na)

n(Na) = n(NaN3) = 1.0 mol


The molar mass of Na is 23 g/mol

Hence,

(1.0 mol Na) × (23 g Na / 1 mol Na) = 23 g Na


Answer: 23 grams of sodium (Na) will result.